OpenStudy (anonymous):

Simplify the following expression: 1/ (\sqrt[3]{x^2}) That is a cube root with the argument x^2 in the denominator. ATTEMPT: Is it not: 1/ (x^2*\sqrt[3]{x^2}) ?

OpenStudy (stormfire1):

\[1/\sqrt[3]{x ^{2}}=1/((x^2)^{1/3})=1/x ^{2/3}\]

OpenStudy (stormfire1):

or maybe more simply: \[x ^{-2/3}\]